# Holomorphic bootstrap for RCFT: signs and bounds for quasi-characters

### v1

@@ 1 Introduction
[…]
@@
The construction of admissible characters can be thought of as the solution to a modular bootstrap problem. They satisfy the basic requirements that lead to a candidate partition function with modular invariance and having a plausible physical interpretation. It has become increasingly clear, however, that the space of actual CFT is far smaller than the space of admissible characters. This fact has long been known in rank-1 (meromorphic) CFTs [13] and in recent years has been explicitly demonstrated for rank $>1$ as well [14, 1, 15, 16]. Thus, the existence of a CFT for a given set of admissible characters is a separate (but important) question. For progress, see [17, 18, 19, 20, 16, 21, 22, 23]. However this question will not be addressed in the present work.
@@ 2.1 Review of quasi-characters
when acting on a weight-$w$ modular function and $E_{2}(\tau)$ is the second Eisenstein series, which transforms as a connection on moduli space. The function $\phi_{2},\phi_{4}$ are modular of weight 2 and 4 respectively under SL(2,Z). In general they are meromorphic and can have ${\textstyle\frac{\ell}{6}}$ poles in $\tau$, where the fractional number comes from the singular points in torus moduli space (see [1] for details) and $\ell$ is an integer called the Wronskian index. The solutions $\chi$ of the equation are of course holomorphic in $\tau$, except at $\tau\to i\infty$.
@@ 2.2 Explicit computations of low-lying coefficients
In this sub-section, as examples, we tabulate the first few coefficients for the ${\mathsf{A}}_{1,1}$ series for $0\leq n\leq 8$ and $-3\leq M\leq 3$.
{adjustwidth}
-35pt0pt $M$ Identity character Non-identity character 0 $1+3q+4q^{2}+7q^{3}+13q^{4}+19q^{5}+29q^{6}+43q^{7}+62q^{8}+{\cal O}(q^{9})$ $1+q+3q^{2}+4q^{3}+7q^{4}+10q^{5}+17q^{6}+23q^{7}+35q^{8}+{\cal O}(q^{9})$ 1 $1-245q+142640q^{2}+18615395q^{3}+837384535q^{4}+21412578125q^{5}+379389640345q^{6}+5165089068645q^{7}+57498950829715q^{8}+{\cal O}(q^{9})$ $1105+101065q+3838295q^{2}+88358360q^{3}+1454696521q^{4}+18742858160q^{5}+199800669415q^{6}+1829051591175q^{7}+14763322626790q^{8}+{\cal O}(q^{9})$ 2 $13-4361q+1024492q^{2}-284433485q^{3}+296843797565q^{4}+84306237909803q^{5}+8867059968079425q^{6}+534104386666020723q^{7}+21861967373053966060q^{8}+{\cal O}(q^{9})$ $10778285+2203389405q+196852441583q^{2}+10638927572924q^{3}+401976188060447q^{4}+11547256546262990q^{5}+266410265073401785q^{6}+5130076445059897723q^{7}+84818281019706292425q^{8}+{\cal O}(q^{9})$ 3 $119-53363q+14459256q^{2}-3364790387q^{3}+842188593869q^{4}-303881533638137q^{5}+461207383305660887q^{6}+203501875932273013375q^{7}+34505840401212977669887q^{8}+{\cal O}(q^{9})$ $72679207559+23350075063007q+3383269741642441q^{2}+300871659609546368q^{3}+18776600055902020059q^{4}+888536031807642974868q^{5}+33555109821077459224757q^{6}+1048799975659078891924625q^{7}+27882387156891739840510632q^{8}+{\cal O}(q^{9})$ $-1$ $209+13547q+360088q^{2}+5912219q^{3}+70940547q^{4}+679136057q^{5}+5469901089q^{6}+38382994913q^{7}+240478794445q^{8}+{\cal O}(q^{9})$ $1-247q-86241q^{2}-4182736q^{3}-96220123q^{4}-1444316844q^{5}-16273561845q^{6}-148598558649q^{7}-1153094272636q^{8}-{\cal O}(q^{9})$ $-2$ $1747563+306927249q+23306923812q^{2}+1067344426501q^{3}+34209156782563q^{4}+836213984046341q^{5}+16482880547977319q^{6}+272348444677096285q^{7}+3880481233093420304q^{8}+{\cal O}(q^{9})$ $11-3397q+796489q^{2}-286457228q^{3}-233881376239q^{4}-31524787414118q^{5}-2067774861752617q^{6}-85699330454705235q^{7}-2550099348343571677q^{8}-{\cal O}(q^{9})$ $-3$ $2228934627+650601251681q+85092063968768q^{2}+6806495569359481q^{3}+381489267486322253q^{4}+16209645539099609007q^{5}+550012236476163613755q^{6}+15464514647907602118167q^{7}+370388802746111712875333q^{8}+{\cal O}(q^{9})$ $19-7973q+2069429q^{2}-480847208q^{3}+130883258691q^{4}-66753637072728q^{5}-86276737034684531q^{6}-19552263390203208491q^{7}-2220250925482531884330q^{8}-{\cal O}(q^{9})$
@@ 2.4 Modular transformations
In the first two lines, the given $S$-matrix is real, symmetric and unitary (hence it squares to 1). In the remaining two lines it is real but neither symmetric nor unitary, but still squares to 1. In the latter two cases it is not the true $S$-matrix. As explained in [3], in these cases one has more than two primary fields (hence they correspond to rank $>2$ MTC). In fact the ${\mathsf{A}}_{2},{\mathsf{E}}_{6}$ series have three primaries, with two of them corresponding to complex conjugates of each other and thus having the same character, while the ${\mathsf{D}}_{4}$ series has four primaries of which three (the $8_{\rm v},8_{\rm c},8_{\rm s}$ representations of ${\mathsf{D}}_{4}$) are related by triality and thus have the same character. Thus the true $S$-matrices should have rank 3 and 4 respectively. The ones listed above then follow by “collapsing” the rows and columns corresponding to the same characters. The true $S$-matrices for these cases were constructed in [3]. They coincide with the known $S$-matrices for the corresponding MTC [7], however we do not need them as the collapsed matrices are sufficient for our purpose here.
@@ 2.5 Combining quasi-characters
[…]
@@ 3.2 $c\gg n$
Empirically, as seen in Table 2, the identity character for $c=24M+j>0$ has exactly $M$ negative coefficients, namely $a_{1},a_{3},\cdots,a_{2M-1}$ while the non-identity character is completely positive. On the other hand for $c=24M+j<0$, the identity character is completely positive while the non-identity has $|M|-1$ positive coefficients namely $a_{2},a_{4},\cdots a_{2|M|-2}$, always assuming that we have chosen $a_{0}>0$. Thus in both cases, precisely at $n=2|M|$ there is a crossover from alternating behaviour to a fixed sign, which takes place in the identity character for $M>0$ and in the non-identity character for $M<0$. We now prove part of the above result: namely that for $n\ll c$, the alternating signs indeed arise as seen above. Later we will examine the crossover region. In the following discussion we set $a_{0}=1$.
@@ 4.1 Recursion relation in the limit
The cases $n=2M,M>0$ for the identity character, and $n=-2M,M<0$ for the non-identity, must be considered separately – here the expression Eq. (4.1) is not valid, and must be replaced by:
@@
$\begin{split}f_{M}(n,i)<0\hbox{ for: }&M>0\hbox{ and }n\leq 2M\qquad\hbox{(identity)}\\ &M<0\hbox{ and }n<2|M|\quad\hbox{(non-identity)}\\ f_{M}(n,i)>0\hbox{ for: }&M<0\quad\hbox{(identity)}\\ &M>0\quad\hbox{(non-identity)}\end{split}$ (4.3)
@@
Next, the ordered partition $(2.1^{n-2})$ corresponds to $p=n-1$ and we have $i_{1}=2;i_{2},i_{3},\cdots,i_{n-1}=1$. The associated term is:
@@
Note that $g_{M}(k,1)=1$. Intuitively, the combinatorial meaning of each $g_{M}(k,m)$ is “start with a chain of $n$ symbols each equal to 1, then at the $k$th position in the chain, delete the next $m$ factors of 1 and replace them by a single factor of $m$”.
@@ 4.2 A crude approximation
It remains to consider $n=2M$ as well as $n>2M$. We saw above that at $n=2M$, the term $n-2M$ in the denominator of Eq. (2.15) vanishes and therefore we must use Eq. (4.2) instead. Thus we have:
@@
To check the validity of our approximations, let us estimate the growth of both $a_{2|M|}$ and $P_{M}(2|M|)$ with $|M|$. First let $M>0$. Extrapolating from explicit computations  [^11] The computation was done for the ${\mathsf{A}}_{1}$ family, i.e. $j=1$ in Eq. (2.9), but is expected to give similar answers for all families, since $j\sim{\cal O}(1)$ as $81\leq j<$ 8., we find that the value of $a_{2M}$ up to $M=1800$ (central charge $c=19200$) is consistent with exponential growth:
@@
Next consider $M<0$. We have already confirmed that $a_{n}$ is positive in this case. As noted above, the upper and lower bounds coming from the Rademacher expansion (Eq. (4.27)) and Eq. (3.5) are the same as for positive $M$. On the other hand, explicit calculation up to around $M=-1800$ tells us that:
@@ 4.3 A better approximation
For large $m$ this tends to $-\frac{1}{240}\frac{\sigma_{3}(m+1)}{\sigma_{3}(m)}$. Since the generalised divisor function $\sigma_{3}(m)$ grows as $m^{3}$, this ratio tends to $-\frac{1}{240}\frac{(m+1)^{3}}{m^{3}}\sim-\frac{1}{240}$ for large $m$, and the series $\sum_{m=2}^{2M}G_{M}(m)$ is convergent. Moreover it is well-approximated by:
@@
The calculated value of the same ratio is in fact found to settle down to $-0.0104$ for sufficiently large $M$ (the error arises from the fact that we neglected the first term in the numerator of Eq. (2.15) and also neglected some small terms along the way). This is a useful confirmation of our estimates in Eq. (4.1). It disagrees with $-\frac{1}{240}$ by a factor $\sim 2.4$, but that was expected due to $m$ being small. This discrepancy will only affect the first few terms in Eq. (4.33) and will not significantly affect Eq. (4.35).
@@
As for the magnitude, the exponential factor leads to a small correction to the “crude” approximation based only on Eq. (4.26). Using that equation in Eq. (4.42) we now get:
@@ 5.2 Bounds on the functions $f_{M}(n,i)$
Next we find bounds on $N_{n,i}$. Expanding $A$ in Eq. (5.2), we see that:
@@
from which:
@@ 5.3 Identity character
$M>0$ case
Now we will inductively show the following: in the range $1\leq n\leq 2M$, $a_{n}<0$ for odd $n$, $a_{n}>0$ for even $n$; and in the range $2M<n<n_{1}\simeq 60M^{2}+\mathcal{O}(M)$, $a_{n}>0$. For us it will be sufficient to consider the range $2M<n\leq 50M^{2}$.
@@
Using Eq. (5.17), we get:
$|f_{M}(n,1)|=\frac{N_{n,1}}{|D_{n}|}>\frac{8(30M+7)}{M+1}\geq 148$ (5.21)
Now we will show that there exists a number $R$ satisfying $1<R\leq 148$ – independent of $n$ – such that $b_{n}\geq Rb_{n-1}$ for all $n\leq 2M$. This in turn will imply that all the $b_{n}$ are positive, which is our goal. The result is already true for $b_{1}$, since $b_{1}=|f_{M}(n,1)|b_{0}\geq 148\,b_{0}$ by Eq. (5.21). Next, fix $n$ and assume the result to be true for all $k<n$. Then we will prove it holds for $k=n$.
@@
Now if we can prove $\alpha(R)\gtrsim R$ then we are done. This in turn will hold if there is a real solution to $\alpha(R)=R$ with $\alpha(R)$ given by Eq. (5.26). Numerically we find there are three real solutions, $\simeq 0.33,53.35,97.93$, with $\alpha(R)>R$ between the latter two. Thus for $R$ anywhere in the range $9754<R<$ 97, the desired inequality holds.
@@
Now let us try to find a bound for $f_{M}(n,i)$ in the current range of interest, $2M<n\leq 50M^{2}$. Following the same procedure used to derive Eq. (5.13), we get:
$\frac{f_{M}(n,i)}{f_{M}(n,1)}$ $=\frac{N_{n,i}}{N_{n,1}}$ (5.31)
$\leq\frac{(291M^{2}+243M+49)i^{3}+4(M+1)i}{40M^{2}+64M+6}$ (5.32)
$<\frac{291M^{2}+247M+53}{40M^{2}+64M+6}i^{3},$
$<8i^{3},$ (5.33)
To get the last inequality we used the fact that $\frac{291M^{2}+247M+53}{40M^{2}+64M+6}$ is an increasing function of $M$ that tends to 7.28 as $M\to\infty$.
@@
since: $01-\sum_{i=2}^{\infty}\frac{8i^{3}}{R^{i-1}}>$ 0 for $R\in(68,97)$. This range is contained in the original one $R\in(54,97)$. So the base case is done: $a_{2M+1}>0$. Now our induction hypothesis is $a_{n}>0$ for $2M+1<n<50M^{2}$, and we want to show that $a_{50M^{2}}>0$.
@@
where, in the second last step, we have used the fact that: $|a_{2M-2k-1}|\leq\frac{|a_{2M-2k}|}{R}$ for all $10\leq k\leq M-$ 1 and $R\in(68,97)$ – which was proven above. Now if we can show that the term in the square bracket above is positive then that would imply that the whole RHS of Eq. (5.35) is positive leading to $a_{n}>0$ for $n=50M^{2}$. So, let us show this next.
Note that for all $i\equiv n-2M+2k$ with $10\leq k\leq M-$ 1 we have $50M^{2}\geq i\geq 50M^{2}-2M>48M^{2}$. From the definitions in Eq. (5.2) and using similar procedure as outlined in Sec. 5.2 we get the following bounds:
@@
for all $10\leq k\leq M-$ 1. Hence, for the RHS of Eq. (5.3) to be positive we need to satisfy $00.782-\frac{1}{R}>$ 0 implying $R\gtrsim 1.23$ which is true since we had picked some $R\in(54,97)$. Hence, we are done.
$M<0$ case
@@
where we have used that $a_{k}>0$ for all $20\leq k\leq n-$ 2. Hence, proved.

### v3

@@ 1 Introduction
$\chi_{i}\left(\frac{a\tau+b}{c\tau+d}\right)=\sum_{j}\varrho_{ij}\begin{pmatrix}a&b\\ c&d\end{pmatrix}\chi_{j},\quad\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\hbox{SL(2,Z)}$ (1.2)
@@
The construction of admissible characters can be thought of as the solution to a modular bootstrap problem. They satisfy the basic requirements that lead to a candidate partition function with modular invariance and having a plausible physical interpretation. It has become increasingly clear, however, that the space of actual CFT is far smaller than the space of admissible characters. This fact has long been known in rank-1 (meromorphic) CFTs [13] and in recent years has been explicitly demonstrated for rank >1>1 as well [14, 1, 15, 16]. Thus, the existence of a CFT for a given set of admissible characters is a separate (but important) question. For progress, see [17, 18, 19, 20, 16, 21, 22, 23]. However this question will not be addressed in the present work.
@@ 2.1 Review of quasi-characters
when acting on a weight-$w$ modular function and $E_{2}(\tau)$ is the second Eisenstein series, which transforms as a connection on moduli space. The function $\phi_{2},\phi_{4}$ are modular of weight 2 and 4 respectively under SL(2,Z). In general they are meromorphic and can have $\frac{\ell}{6}$} poles in $\tau$, where the fractional number comes from the singular points in torus moduli space (see [1] for details) and $\ell$ is an integer called the Wronskian index. The solutions $\chi$ of the equation are of course holomorphic in $\tau$, except at $\tau\to i\infty$.
@@ 2.2 Explicit computations of low-lying coefficients
In this sub-section, as examples, we tabulate the first few coefficients for the ${\mathsf{A}}_{1,1}$ series for $0\leq n\leq 8$ and $3\leq M\leq 3$.
$M$ Identity character Non-identity character
0 $1+3q+4q^{2}+7q^{3}+13q^{4}+19q^{5}+29q^{6}+43q^{7}+62q^{8}+{\cal O}(q^{9})$ $1+q+3q^{2}+4q^{3}+7q^{4}+10q^{5}+17q^{6}+23q^{7}+35q^{8}+{\cal O}(q^{9})$
[…]
2 $13-4361q+1024492q^{2}-284433485q^{3}+296843797565q^{4}+84306237909803q^{5}+8867059968079425q^{6}+534104386666020723q^{7}+21861967373053966060q^{8}+{\cal O}(q^{9})$ $10778285+2203389405q+196852441583q^{2}+10638927572924q^{3}+401976188060447q^{4}+11547256546262990q^{5}+266410265073401785q^{6}+5130076445059897723q^{7}+84818281019706292425q^{8}+{\cal O}(q^{9})$
3 $119-53363q+14459256q^{2}-3364790387q^{3}+842188593869q^{4}-303881533638137q^{5}+461207383305660887q^{6}+203501875932273013375q^{7}+34505840401212977669887q^{8}+{\cal O}(q^{9})$ $72679207559+23350075063007q+3383269741642441q^{2}+300871659609546368q^{3}+18776600055902020059q^{4}+888536031807642974868q^{5}+33555109821077459224757q^{6}+1048799975659078891924625q^{7}+27882387156891739840510632q^{8}+{\cal O}(q^{9})$
1-1 $209+13547q+360088q^{2}+5912219q^{3}+70940547q^{4}+679136057q^{5}+5469901089q^{6}+38382994913q^{7}+240478794445q^{8}+{\cal O}(q^{9})$ $1-247q-86241q^{2}-4182736q^{3}-96220123q^{4}-1444316844q^{5}-16273561845q^{6}-148598558649q^{7}-1153094272636q^{8}-{\cal O}(q^{9})$
[…]
3-3 $2228934627+650601251681q+85092063968768q^{2}+6806495569359481q^{3}+381489267486322253q^{4}+16209645539099609007q^{5}+550012236476163613755q^{6}+15464514647907602118167q^{7}+370388802746111712875333q^{8}+{\cal O}(q^{9})$ $19-7973q+2069429q^{2}-480847208q^{3}+130883258691q^{4}-66753637072728q^{5}-86276737034684531q^{6}-19552263390203208491q^{7}-2220250925482531884330q^{8}-{\cal O}(q^{9})$
@@ 2.4 Modular transformations
In the first two lines, the given $S$-matrix is real, symmetric and unitary (hence it squares to 1). In the remaining two lines it is real but neither symmetric nor unitary, but still squares to 1. In the latter two cases it is not the true $S$-matrix. As explained in [3], in these cases one has more than two primary fields (hence they correspond to rank >2>2 MTC). In fact the ${\mathsf{A}}_{2},{\mathsf{E}}_{6}$ series have three primaries, with two of them corresponding to complex conjugates of each other and thus having the same character, while the ${\mathsf{D}}_{4}$ series has four primaries of which three (the $8_{\rm v},8_{\rm c},8_{\rm s}$ representations of ${\mathsf{D}}_{4}$) are related by triality and thus have the same character. Thus the true $S$-matrices should have rank 3 and 4 respectively. The ones listed above then follow by “collapsing” the rows and columns corresponding to the same characters. The true $S$-matrices for these cases were constructed in [3]. They coincide with the known $S$-matrices for the corresponding MTC [7], however we do not need them as the collapsed matrices are sufficient for our purpose here.
@@ 2.5 Combining quasi-characters
Notice that if $M_{\rm max}+M_{\rm min}+\frac{j+2}{12}<0$ then $h$ is negative and we necessarily have a non-unitary theory. Now Eq. (2.9) implies that $\frac{1}{4}\leq\frac{j+2}{12}\leq\frac{3}{4}$, so for any positive $M_{\rm max}$ the largest allowed number of terms in the sum consistent with unitarity is attained by choosing $M_{\rm min}=-M_{\rm max}$. Also the only case for which $h<1$ is $M_{\rm max}=M_{\rm min}=0$. With two characters and $h$ non-integer, the only primary in the full non-chiral CFT (if one exists) has dimensions $(h,h)$. Hence all cases $M_{\rm max}\neq 0$, or more precisely the CFTs among them, correspond to what have been called “perfect metals” in [27] – such CFTs have no relevant deformations.
@@ 3.2 $c\gg n$
Empirically, as seen in Table 2, the identity character for $c=24M+j>0$ has exactly $M$ negative coefficients, namely $a_{1},a_{3},\cdots,a_{2M-1}$ while the non-identity character is completely positive. On the other hand for $c=24M+j<0$, the identity character is completely positive while the non-identity has |M|1|M|-1 positive coefficients namely $a_{2},a_{4},\cdots a_{2|M|-2}$, always assuming that we have chosen $a_{0}>0$. Thus in both cases, precisely at $n=2|M|$ there is a crossover from alternating behaviour to a fixed sign, which takes place in the identity character for $M>0$ and in the non-identity character for $M<0$. We now prove part of the above result: namely that for $n\ll c$, the alternating signs indeed arise as seen above. Later we will examine the crossover region. In the following discussion we set $a_{0}=1$.
@@ 4.1 Recursion relation in the limit
The cases $n=2M,M>0$ for the identity character, and n=2M,M<0n=-2M,M<0 for the non-identity, must be considered separately – here the expression Eq. (4.1) is not valid, and must be replaced by:
@@
[…]
@@
Next, the ordered partition $(2.1^{n-2})$ corresponds to p=n1p=n-1 and we have $i_{1}=2;i_{2},i_{3},\cdots,i_{n-1}=1$. The associated term is:
@@
Note that $g_{M}(k,1)=1$. Intuitively, the combinatorial meaning of each $g_{M}(k,m)$ is “start with a chain of $n$ symbols each equal to 1, then at the $k$th position in the chain, delete the next $m$ factors of 1 and replace them by a single factor of mm”.
@@ 4.2 A crude approximation
It remains to consider $n=2M$ as well as $n>2M$. We saw above that at $n=2M$, the term n2Mn-2M in the denominator of Eq. (2.15) vanishes and therefore we must use Eq. (4.2) instead. Thus we have:
@@
To check the validity of our approximations, let us estimate the growth of both $a_{2|M|}$ and $P_{M}(2|M|)$ with $|M|$. First let $M>0$. Extrapolating from explicit computations  [^11] The computation was done for the ${\mathsf{A}}_{1}$ family, i.e. $j=1$ in Eq. (2.9), but is expected to give similar answers for all families, since $j\sim{\cal O}(1)$ as $1\leq j<8$., we find that the value of $a_{2M}$ up to $M=1800$ (central charge $c=19200$) is consistent with exponential growth:
@@
Next consider $M<0$. We have already confirmed that $a_{n}$ is positive in this case. As noted above, the upper and lower bounds coming from the Rademacher expansion (Eq. (4.27)) and Eq. (3.5) are the same as for positive $M$. On the other hand, explicit calculation up to around M=1800M=-1800 tells us that:
@@ 4.3 A better approximation
For large $m$ this tends to $\frac{1}{240}\frac{\sigma_{3}(m+1)}{\sigma_{3}(m)}$. Since the generalised divisor function $\sigma_{3}(m)$ grows as $m^{3}$, this ratio tends to $\frac{1}{240}\frac{(m+1)^{3}}{m^{3}}\sim-\frac{1}{240}$ for large $m$, and the series $\sum_{m=2}^{2M}G_{M}(m)$ is convergent. Moreover it is well-approximated by:
@@
The calculated value of the same ratio is in fact found to settle down to 0.0104-0.0104 for sufficiently large $M$ (the error arises from the fact that we neglected the first term in the numerator of Eq. (2.15) and also neglected some small terms along the way). This is a useful confirmation of our estimates in Eq. (4.1). It disagrees with $\frac{1}{240}$ by a factor $\sim 2.4$, but that was expected due to $m$ being small. This discrepancy will only affect the first few terms in Eq. (4.33) and will not significantly affect Eq. (4.35).
@@
Now for the magnitude of $a_{2M}$, the exponential factor leads to a small correction to the “crude” approximation based only on Eq. (4.26). Using that equation in Eq. (4.42) we now get:
@@
In the next section, we will provide an inductive argument to show that the sign of $a_{n}$ is the same as sign of $(-1)^{n}$. Using this result in appendix B we will show that the exact expression for $a_{n}$ as in Eq. (4.7) can be written as an exponential function of $M$ and a real positive parameter $\beta$ which we define in that appendix. This would imply that for $1\leq n\leq 2M$, $a_{n}$ as given in Eq. (4.7) nicely exponentiates!
@@ 5.2 Bounds on the functions $f_{M}(n,i)$
Now let us take $M>0$. We find bounds on $N_{n,i}$. Expanding $A$ in Eq. (5.2), we see that:
@@
from which we find (with $M>0$):
@@
Above will be precisely the two regimes of $n$ that we will be focusing on during the inductive proof for the identity character  [^15] For simplicity, we will take the upper bound in the second regime of $n$ to be $50M^{2}$ instead..
@@
$\mathbf{1}\leq\bm{n}\leq\mathbf{2}\bm{M}$
Here, we will restrict for the first range: $1\leq n\leq 2M$. Let us put $n=1$ in Eq. (5.17) to get:
$|f_{M}(n,1)|=\frac{N_{n,1}}{|D_{n}|}>\frac{8(30M+7)}{M+1}\geq 148,$ (5.18)
Eqs. (5.18) and (5.19) are the kind of inequalities that we will use in the inductive proof later.
$\bm{2M}<\bm{n}\leq\bm{50M^{2}}$
Now let us try to find a bound for $f_{M}(n,i)$ in the second range of interest, $2M<n\leq 50M^{2}$. Following the same procedure used to derive Eq. (5.13), we get:
$\frac{f_{M}(n,i)}{f_{M}(n,1)}$ $=\frac{N_{n,i}}{N_{n,1}}$ (5.21)
$\leq\frac{(291M^{2}+243M+49)i^{3}+4(M+1)i}{40M^{2}+64M+6}$ (5.22)
$<\frac{291M^{2}+247M+53}{40M^{2}+64M+6}i^{3},$
$<8i^{3},$ (5.23)
To obtain the last inequality, we used the fact that $\frac{291M^{2}+247M+53}{40M^{2}+64M+6}$ is an increasing function of $M$ that tends to 7.28 as $M\to\infty$.
@@ 5.3 Identity character
$\bm{M}>\bm{0}$ case
In this section, we will inductively show the following: in the range $1\leq n\leq 2M$, $a_{n}<0$ for odd $n$, $a_{n}>0$ for even $n$; and in the range $2M<n<n_{1}\simeq 60M^{2}+\mathcal{O}(M)$, $a_{n}>0$. For us it will be sufficient to consider the range $2M<n\leq 50M^{2}$.
@@
In order to show this, we will first show that there exists a number $R$ satisfying $1<R\leq 148$ – independent of $n$ – such that $b_{n}\geq Rb_{n-1}$ for all $n\leq 2M$. This in turn will imply that all the $b_{n}$ are positive, which is our goal. The result is already true for $b_{1}$, since $b_{1}=|f_{M}(n,1)|b_{0}\geq 148\,b_{0}$ by Eq. (5.18). Next, fix $n$ and assume the result to be true for all $k<n$. Then we will prove it holds for $k=n$.
@@
Now if we can prove $\alpha(R)\gtrsim R$ then we are done. This in turn will hold if there is a real solution to $\alpha(R)=R$ with $\alpha(R)$ given by Eq. (5.30). Numerically we find there are three real solutions, $\simeq 0.33,53.35,97.93$, with $\alpha(R)>R$ between the latter two. Thus for $R$ anywhere in the range $54<R<97$, the desired inequality holds.
@@
since: $1-\sum_{i=2}^{\infty}\frac{8i^{3}}{R^{i-1}}>0$ for $R\in(68,97)$. This range is contained in the original one $R\in(54,97)$. So the base case is done: $a_{2M+1}>0$. Now our induction hypothesis is $a_{n}>0$ for $2M+1<n<50M^{2}$, and we want to show that $a_{50M^{2}}>0$.
@@
where, in the second last step, we have used the fact that: $|a_{2M-2k-1}|\leq\frac{|a_{2M-2k}|}{R}$ for all $0\leq k\leq M-1$ and $R\in(68,97)$ – which was proven above. Now if we can show that the term in the square bracket above is positive then that would imply that the whole RHS of Eq. (5.35) is positive leading to $a_{n}>0$ for $n=50M^{2}$. So, let us show this next.
Note that for all $i\equiv n-2M+2k$ with $0\leq k\leq M-1$ we have $50M^{2}\geq i\geq 50M^{2}-2M>48M^{2}$. From the definitions in Eq. (5.2) and using similar procedure as outlined in Sec. 5.2 we get the following bounds:
@@
for all $0\leq k\leq M-1$. Hence, for the RHS of Eq. (5.3) to be positive we need to satisfy $0.782-\frac{1}{R}>0$ implying $R\gtrsim 1.23$ which is true since we had picked some $R\in(54,97)$. Hence, we are done.
$\bm{M}<\bm{0}$ case
@@
where we have used that $a_{k}>0$ for all $0\leq k\leq n-2$. Hence, proved.